# DPZV: Elevating the Tradeoff between Privacy and Utility in Zeroth-Order Vertical Federated Learning

### v2

@@ 2. $(ii)$
[…]
This is the standard result of zeroth-order optimization. The proof of the Lemma is given by [28]
We also find the following lemmas useful in the proof:
Lemma A.7.
@@
This lemma follows exactly from Lemma C.1. in [38]. ∎
Lemma A.8.
Let $Q$ be the event that clipping happened for a sample $\xi$, $d$ be the model dimension, and $L,\ell$ be the Lipschitz and smooth constant as defined in A.1 and A.2. For $\forall C_{0}>0$, if the clipping threshold $C$ follows $C\geq C_{0}+L\lambda d/2$, we have the following upper bound for the probability of clipping:
@@
Therefore, by Lemma A.7 and A.1, we obtain
@@
Lemma A.9 (Expectation and Variance of Clipped Zeroth-order Gradient Estimator).
Recall that $\breve{g}^{t}_{m,i}({\bm{w}}^{t})$ is defined as the clipped zeroth-order gradient estimator assuming no communication delay, random perturbation ${\bm{u}}$ is defined in Lemma A.7, and event $Q$ is defined in Lemma A.8. We have the following properties:
@@
where in the first step we applied the Law of Total Expectation, and in the last step we used the property $(i)$ in Lemma A.6 and $(i)$ in Lemma A.7.
@@
where 1)1)1 ) is by the property of zeroth-order gradient estimator (15) and 2)2)2 ) follows from the bounded gradient assumption(Assumption A.3). ∎
Lemma A.10 (Bounds on the variance of DP Zeroth-order Gradient Estimator).
Let $\breve{G}^{t}_{m}({\bm{w}}^{t})$ be the Differential Private Zeroth-order Gradient without delay. Under the same condition as Lemma A.6, we can bound the variance of $\breve{G}^{t}_{m}({\bm{w}}^{t})$ in expectation, with the expectation taken on random direction ${\bm{u}}$, DP noise $z$, and clipping event $Q$:
@@
Finally, for analyzing the delayed information in model parameter, we define the following Lyapunov function:
with $\gamma_{i}$ to be determined later.
Appendix B Proof of Convergence Theorem 5.4
We first state the full version of the main convergence Theorem 5.4.
Theorem B.1.
Under A.1 to A.4, and assume the client delay $\tau_{m}$ is uniformly upper bounded by $\tau$. If we denote $q_{*}=\min_{m}q_{m}$, $d_{*}=\max_{m}d_{m}$, and let $\eta_{0}=\eta_{m}=\eta\leq\min\{\frac{1}{\sqrt{Td_{*}}},\frac{B}{4L(B+8d_{0})+8\gamma_{1}(2d_{m}+B)}\}$, $\lambda\leq\frac{1}{Ld\sqrt{T}}$, we have the following theorem:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{{\bm{w}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]\leq$ $\frac{1}{1-\Xi}\left\{\frac{8\sqrt{d_{*}}}{q_{*}T^{1/2}}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{2d_{*}}{q_{*}T}+\frac{8\sqrt{d_{*}}}{q_{*}LdT^{1/2}}+\frac{16(4L+2\gamma_{1})\sqrt{d_{*}}\sigma_{s}^{2}}{q_{*}BT^{1/2}}\right.$
$+$ $\left.\frac{4\left((4-B)L+(B+2)\gamma_{1}\right)\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})\Xi C^{2}\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})\sigma_{dp}^{2}d_{m}}{q_{*}T^{1/2}\sqrt{d_{*}}}\right\},$ (23)
where $\Xi=2\sqrt{2\pi}\exp(-\frac{C_{0}^{2}}{8\ell^{2}}).$
Proof.
We start from the result of an intermediate Lemma D.2, which quantifies the descent of the Lyapunov function, and we relegate the proof to Appendix D.
$\mathbb{E}\left[V^{t+1}-V^{t}\right]\leq-\frac{1-P}{8}\min_{m}\{q_{m}\eta_{m}\}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{{\bm{w}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]+{\mathcal{A}}_{1}+{\mathcal{A}}_{2}$ (24)
Note that ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ are constant terms defined in (35) and (39).
We first re-arranging the terms in (24), and take average over $10,1,\dots,T-$ 1:
where the second inequality follows from the definition of $V^{t}$ in (22).
$+\frac{1}{\alpha}\sum_{m=0}^{M}q_{m}\eta_{m}^{2}L\left(\frac{4}{B}d_{m}\sigma_{s}^{2}+\frac{(4-B)L^{2}}{4B}\lambda^{2}d_{m}^{3}+\frac{2P}{B}C^{2}d_{m}+2\sigma_{dp}^{2}d_{m}\right)$
$+\frac{1}{\alpha}L\lambda^{2}d+\frac{1}{\alpha}\sum_{m=1}^{M}q_{m}\eta_{m}^{2}\gamma_{1}\left(\frac{L^{2}}{4}\lambda^{2}d_{m}^{3}+\frac{2}{B}d_{m}\sigma_{s}^{2}+\frac{L^{2}}{2B}\lambda^{2}d_{m}^{3}+\frac{P}{B}C^{2}d_{m}+\sigma_{dp}^{2}d_{m}\right)$
$\leq$ $\frac{8}{q_{*}\eta T}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{2d_{*}^{3}/d^{2}}{q_{*}T}+\frac{16\eta L}{q_{*}}\left(\frac{4}{B}d_{*}\sigma_{s}^{2}+\frac{(4-B)d_{*}^{3}/d^{2}}{4BT}+\frac{2P}{B}C^{2}d_{*}+2\sigma_{dp}^{2}d_{*}\right)$
$+$ $\frac{8}{q_{*}Ld\eta T}+\frac{16\eta\gamma_{1}}{q_{*}}\left(\frac{2}{B}d_{*}\sigma_{s}^{2}+\frac{(B+2)d_{*}^{3}/d^{2}}{4BT}+\frac{P}{B}C^{2}d_{*}+\sigma_{dp}^{2}d_{m}\right)$
$\leq$ $\frac{8}{q_{*}\eta T}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{2d_{*}}{q_{*}T}+\frac{8}{q_{*}Ld\eta T}+\frac{16\eta(4L+2\gamma_{1})d_{*}\sigma_{s}^{2}}{q_{*}B}+\frac{4\eta\left((4-B)L+(B+2)\gamma_{1}\right)d_{*}}{q_{*}BT}$
where in the last step, we use the fact that $d>d_{*}$.
If we choose $\eta=\frac{1}{\sqrt{Td_{*}}}$, we can get the convergence rate:
$\leq$ $\frac{1}{1-P}\left\{\frac{8\sqrt{d_{*}}}{q_{*}T^{1/2}}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{2d_{*}}{q_{*}T}+\frac{8\sqrt{d_{*}}}{q_{*}LdT^{1/2}}+\frac{16(4L+2\gamma_{1})\sqrt{d_{*}}\sigma_{s}^{2}}{q_{*}BT^{1/2}}\right.$
$+$ $\left.\frac{4\left((4-B)L+(B+2)\gamma_{1}\right)\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})PC^{2}\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})\sigma_{dp}^{2}\sqrt{d_{*}}}{q_{*}T^{1/2}}\right\}$
$\leq$ $\frac{1}{1-\Xi}\left\{\frac{8\sqrt{d_{*}}}{q_{*}T^{1/2}}\mathbb{E}\left[F^{0}-F^{T}\right]+\frac{2d_{*}}{q_{*}T}+\frac{8\sqrt{d_{*}}}{q_{*}LdT^{1/2}}+\frac{16(4L+2\gamma_{1})\sqrt{d_{*}}\sigma_{s}^{2}}{q_{*}BT^{1/2}}\right.$
$+$ $\left.\frac{4\left((4-B)L+(B+2)\gamma_{1}\right)\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})\Xi C^{2}\sqrt{d_{*}}}{q_{*}BT^{1/2}}+\frac{16(2L+\gamma_{1})\sigma_{dp}^{2}\sqrt{d_{*}}}{q_{*}T^{1/2}}\right\}$
where the last step follows by Lemma A.8.
We thus conclude that the convergence rate is
${\mathcal{O}}(\frac{d_{*}^{1/2}\mathbb{E}\left[F^{0}-F^{T}\right]+d_{*}^{1/2}\sigma_{s}^{2}+d_{*}^{1/2}\sigma_{dp}^{2}}{T^{1/2}}),$
where constants before terms have been omitted for simplicity. We note that our convergence rate is of the same order compared to centralized ZO optimization under the same smoothness assumption, up to an error term $\sigma_{dp}$ introduced by Differential Privacy. ∎
Appendix C Proof of Differential Privacy Theorem 6.1
In this section, we give rigorous proof for the differential privacy guarantee in Theorem 6.1.
We first introduce the definition of Gaussian differential privacy (GDP) [9] which will be useful in the proof for Theorem 6.1. Compared with traditional DP defined in (4), this notion of privacy provides a much tighter composition theorem.
Definition C.1 (Gaussian Differential Privacy).
$T(M(X),M(X^{\prime}))\geq G_{\mu}$
For all neighboring dataset $X,X^{\prime}\in{\mathcal{X}}^{n}$.
We construct our DP algorithm based on the Gaussian Mechanism. The Gaussian Mechanism of GDP is given by the following theorem.
Theorem C.2 (Gaussian Mechanism for Gaussian Differential Privacy).
The iterative nature of gradient-like algorithms calls for the composition theorem.
Theorem C.3 (Composition of Gaussian Differential Privacy).
The $T$-fold composition of $\mu$-GDP mechanisms is $\sqrt{T}\mu$-GDP
For the ease of comparison, we convert GDP to the common $(\epsilon,\delta)$-DP based on the following lossless conversion:
Theorem C.4 (Conversion from Gaussian Differential Privacy to $(\epsilon,\delta)$-Differential Privacy).
$\delta(\epsilon)=\Phi(-\frac{\epsilon}{\mu}+\frac{\mu}{2})-e^{\epsilon}\Phi(-\frac{\epsilon}{\mu}-\frac{\mu}{2})$
The proof of Theorem C.2-C.4 is given in [9].
We are now ready to present the proof for Theorem 6.1.
Proof of Theorem 6.1
Proof.
First recall the definition of $\Delta_{m}^{t}$ defined in (8):
$r=\frac{B}{D}$
Note that in Theorem 6.1, the standard variance of $z_{m}^{t}$ is given by
By Theorem C.2, the mechanism conducted in (8) satisfies $(\mu/\sqrt{T})$-Gaussian Differential Privacy. Further applying the composition of GDP in Theorem C.3, and by the post-processing [11] of differential privacy, we have that the whole training process of Algorithm 1 is $\mu$-GDP. We complete the proof by converting $\mu$-GDP to $(\epsilon,\delta)$-DP according to Theorem C.4. ∎
Appendix D Intermediate Lemmas
Lemma D.1 (Model Update With Delay).
Under A.1 to A.4, we have the following lemma:
@@ Appendix D Intermediate Lemmas
where in the second step we take expectation on both sides, first w.r.t. the random direction $u$, DP noise $z$, and the clipping event $Q$, then w.r.t the client $m_{k}$.
We bound ${\mathcal{E}}_{1}$ as the following:
@@
=$\overset{1)}{=}$ $-(1-P)\eta_{0}\mathbb{E}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\nabla_{w_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t})\rangle$
@@
where in 1)1)1 ) we use the fact that $\mathbb{E}\left[\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})\right]=0$, in 2)2)2 ) we applied the Cauchy–Schwarz inequality, and 3)3)3 ) follows by the smoothness of $F_{\lambda}$ and the fact that $1-P\leq 1$
For ${\mathcal{E}}_{2}$, we can further bound it based on Assumption A.2:
@@
where in 1)1)1 ) and 2)2)2 ) we applied the Cauchy–Schwarz inequality, and in 3)3)3 ) we substitute (21) in and use the $L$-smoothness of $F_{\lambda}$, and in $(iv)$ we use the fact that $1-P\leq 1$ and let
@@
where in the last inequality we further take expectation w.r.t. client $M$ and combine similar terms.
@@
where 1)1)1 ) and 2)2)2 ) follows from equation (14) and (17) in Lemma A.6 respectively. In 3)3)3 ), we let $q_{0}=1$ and combine similar terms. In 4)4)4 ), we substitute in (32). Lastly, in 5)5)5 ), we denote
@@
Now, recall the definition of the Lyapunov function $V^{t}$:
$V^{t}=F({\bm{w}}^{t})+\sum_{i=1}^{\tau}\gamma_{i}\mbox{$\left\lVert\chi^{t+1-i}-\chi^{t-i}\right\rVert$}^{2}$
We utilize the Lyapunov function to prove the following lemma for eliminating model delay.
Lemma D.2.
Under A.1-A.4, and assume the client delay $\tau_{m}$ is uniformly upper bounded by $\tau$, we have the following lemma:
@@
Before we give the proof of Lemma D.2, we first provide some useful facts that reveal properties of the delayed parameters.
Recall that $\tilde{\chi}^{t}$ denote the delayed parameters on all clients, and $\chi^{t}$ denote the non-delayed version. Let ${\mathcal{F}}_{1}=\mathbb{E}\left[\mbox{$\left\lVert\chi^{t+1}-\chi^{t}\right\rVert$}^{2}\right]$, ${\mathcal{F}}_{2}=\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]$.
@@
=$\overset{3)}{=}$ $-\eta_{0}(1-P)\left(\frac{1}{4}-\frac{\eta_{0}L(B+8d_{0})}{2B}\right)\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]$
@@
Above, we used Lemma D.1 in step $(1)$, substituted in (36) for ${\mathcal{F}}_{1}$ in step $(2)$, substituted in (37) for ${\mathcal{F}}_{2}$ in step (3), and defined
@@
Appendix E Discussion of Differential Privacy under Various Settings
Overview of DPZV’s Privacy Mechanisms.  DPZV protects privacy through two main mechanisms: 1) Zeroth-Order (ZO) Optimization. By eliminating the need to transmit backward gradients, DPZV mitigates conventional gradient-based leakage [16]. Specifically, no unperturbed gradients are ever exchanged among participants. 2) Controllable Differential Privacy (DP). Each participant interacts only through black-box queries and responses augmented with DP-based noise. This randomized protocol further shields sensitive information from reconstruction or inference attacks.
E.1 Protection Against an Honest-but-Curious Threat Model
In an honest-but-curious scenario, adversaries seek private insights (labels or features) without deviating from the protocol. Two common types of attacks are:
Label Inference Attacks.
Such attacks typically exploit exact gradient signs or rely on known model architectures to backtrack label information [13, 21]. Under DPZV, unbiased gradients are never shared, and the server model remains inaccessible to other participants. Clients observe only stochastic ZO outputs with added noise, preventing them from reverse-engineering labels via gradient signals.
Feature Inference Attacks.
Adversaries may attempt feature or data reconstruction [16] by leveraging gradient disclosures or conducting model inversion. DPZV thwarts such efforts by disallowing adaptive queries on the client’s internal parameters; only scalar function evaluations flow to and from the server. Additionally, the injected DP noise obscures potential patterns that attackers might exploit for reconstructing sensitive features.
Moreover, both label and feature inference methods often assume knowledge of the dataset domain or direct model access, which is not given in DPZV’s design. Task details (e.g., label distributions, model layers) are held privately by each party, limiting the adversary’s capacity to launch sophisticated inversion or inference attacks.
E.2 Protection Against Post-Training Attacks
Post-training adversaries typically aim to infer sensitive data from a final, trained model [2]. Since DPZV applies differential privacy throughout training, the final model parameters satisfy rigorous $(\epsilon,\delta)$-DP guarantees. Hence, even if the trained model is released or accessed, the level of noise injected ensures that the adversary cannot reliably distinguish any single individual’s data—reducing vulnerability to membership inference or model-inversion attacks [18]. This formal DP framework remains robust regardless of downstream usage or queries on the finalized model.
In summary, by combining ZO optimization with DP noise injection, DPZV provides comprehensive protection against both honest-but-curious and post-training adversaries across diverse privacy threat scenarios.
Appendix F Memory Cost
Figure 6 compares the GPU memory consumption, with memory values normalized for readability. We compare the memory cost on larger models, where we use ResNet for image classification and BERT for sequence classification. We record the highest memory peak to show the total required memory for each method on training large models. We observe that DPZV requires memory approximately equal to the model size, whereas the first-order method VAFL demands more than twice the model size. Compared to ZOO-VFL, DPZV achieves further memory savings by leveraging MeZO. These results highlight the scalability of DPZV, making it well-suited for deploying large pretrained language models in VFL scenarios.
Appendix G Choice of all Hyperparameters
In this section, we show all the chosen Hyperparameters, including learning rate for the four datasets and ZO and DP parameters in Sec. 7:
Table 2: All learning rate choices
Dataset VAFL ZOO-VFL VFL-CZOFO DPZV (ours)
MNIST $\eta_{0}$ $0.005$ $0.005$ $0.005$ $0.005$
$\eta_{m}$ $0.001$ $5\times 10^{-5}$ $1\times 10^{-7}$ $5\times 10^{-4}$
CIFAR-10 $\eta_{0}$ $0.005$ $0.005$ $0.005$ $0.005$
$\eta_{m}$ $0.001$ $5\times 10^{-5}$ $5\times 10^{-8}$ $1\times 10^{-4}$
ModelNet40 $\eta_{0}$ $1\times 10^{-4}$ $1\times 10^{-5}$ $5\times 10^{-7}$ $1\times 10^{-5}$
$\eta_{m}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $5\times 10^{-7}$ $1\times 10^{-5}$
Amazon Review $\eta_{0}$ $1\times 10^{-5}$ $5\times 10^{-7}$ $1\times 10^{-6}$ $5\times 10^{-7}$
$\eta_{m}$ $1\times 10^{-5}$ $5\times 10^{-7}$ $1\times 10^{-6}$ $5\times 10^{-7}$
Table 3: Other Hyperparameters
Parameter Value Explanation
$C$ $10$ DP clipping threshold
$\lambda$ $0.001$ Scale of perturbation for ZOO
$m$ $0.9$ Momentum parameter
$\delta$ $1\times 10^{-5}$ DP error level
$n$ $5$ Number of perturbation for VFL-CZOFO

### v3

@@ VIII Conclusion and Future Work
Assumption .1 ($\ell$-Lipschitz).
Assumption .2 ($L$-Smooth).
The function $f({\bm{w}};\xi)$ is $L$-Smooth for every $\xi$. Specifically, there exists an $L>0$ for all $m=0,\dots,M$ such that $\mbox{$\left\lVert\nabla_{w_{m}}f({\bm{w}})-\nabla_{w_{m}}f({\bm{w}}^{\prime})\right\rVert$}\leq L\mbox{$\left\lVert{\bm{w}}-{\bm{w}}^{\prime}\right\rVert$}$.
Assumption .3 (Bounded gradient variance).
Assumption .4 (Independent Participation).
The probability of one device participating in one communication round is independent of other devices and satisfies
$\mathbb{P}(\text{device }m\text{ uploading})=q_{m}$
One of the important parts in the proof of Theorem V.4 is to bound the zeroth-order gradient estimator. We first introduce the formal definition of zeroth-order two-point gradient estimator that is used in our algorithm, and prove some technical lemmas that reveal some important properties.
Definition .5.
Lemma .6 (Restatement of Lemma V.6).
  1. $(i)$
  2. $(ii)$
    If $f({\bm{w}})$ is $L$-smooth, we have that
    $\mbox{$\left\lVert\nabla f({\bm{w}})-\nabla f_{\lambda}({\bm{w}})\right\rVert$}\leq\frac{L}{2}\lambda d^{3/2},$ (25)
@@ 2. $(ii)$
$\mbox{$\left\lVert\nabla f_{\lambda}({\bm{w}})\right\rVert$}^{2}\leq 2\mbox{$\left\lVert\nabla f({\bm{w}})\right\rVert$}^{2}+\frac{L^{2}}{2}\lambda^{2}d^{3}$ (28)
$\mbox{$\left\lVert\nabla f({\bm{w}})\right\rVert$}^{2}\leq 2\mbox{$\left\lVert\nabla f_{\lambda}({\bm{w}})\right\rVert$}^{2}+\frac{L^{2}}{2}\lambda^{2}d^{3}$ (29)
This is the standard result of zeroth-order optimization. The proof of the Lemma is given by [nesterov2017random] We also find the following lemmas useful in the proof:
Lemma .7.
@@
This lemma follows exactly from Lemma C.1. in [zhang2024dpzero]. ∎
Lemma .8 (Restatement of Lemma V.7).
Let $Q$ be the event that clipping happened for a sample $\xi$, $d$ be the model dimension, and $L,\ell$ be the Lipschitz and smooth constant as defined in assumption .1 and assumption .2. For $\forall C_{0}>0$, if the clipping threshold $C$ follows $C\geq C_{0}+L\lambda d/2$, we have the following upper bound for the probability of clipping:
@@
Therefore, by Lemma .7 and Assumption .1, we obtain
@@
Lemma .9 (Expectation and Variance of Clipped Zeroth-order Gradient Estimator).
Recall that $\breve{g}^{t}_{m,i}({\bm{w}}^{t})$ is defined as the clipped zeroth-order gradient estimator assuming no communication delay, random perturbation $\bm{u}$} is defined in Lemma .7, and event $Q$ is defined in Lemma V.7. We have the following properties:
@@
where in the first step we applied the Law of Total Expectation, and in the last step we used the property $(i)$ in Lemma V.6 and $(i)$ in Lemma .7.
@@
where $1)$ is by the property of zeroth-order gradient estimator (15) and $2)$ follows from the bounded gradient assumption(Assumption .3). ∎
Lemma .10 (Restatement of Lemma V.8).
Let $\breve{G}^{t}_{m}({\bm{w}}^{t})$ be the Differential Private Zeroth-order Gradient without delay. Under the same condition as Lemma V.6, we can bound the variance of $\breve{G}^{t}_{m}({\bm{w}}^{t})$ in expectation, with the expectation taken on random direction $\bm{u}$}, DP noise $z$, and clipping event $Q$:
@@
-B Proof of Intermediate Lemmas
Lemma .11 (Restatement of Lemma V.9).
Under Assumption .1 to .4, we have the following lemma:
@@
where in the second step we take expectation on both sides, first w.r.t. the random direction $u$, DP noise $z$, and the clipping event $Q$, then w.r.t the device $m_{k}$. We bound ${\mathcal{E}}_{1}$ as the following:
@@
$\overset{1)}{=}$ $-(1-P)\eta_{0}\mathbb{E}\langle\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t}),\nabla_{w_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})-\nabla_{w_{0}}F_{\lambda}({\bm{w}}^{t})\rangle$
@@
where in $1)$ we use the fact that $\mathbb{E}\left[\breve{G}^{t}_{0}(\tilde{{\bm{w}}}^{t})-(1-P)\nabla_{x_{0}}F_{\lambda}(\tilde{{\bm{w}}}^{t})\right]=0$, in $2)$ we applied the Cauchy–Schwarz inequality, and $3)$ follows by the smoothness of $F_{\lambda}$ and the fact that $1-P\leq 1$
For ${\mathcal{E}}_{2}$, we can further bound it based on Assumption .2:
@@
where in $1)$ and $2)$ we applied the Cauchy–Schwarz inequality, and in $3)$ we substitute (17) in and use the $L$-smoothness of $F_{\lambda}$, and in $(iv)$ we use the fact that $1-P\leq 1$ and let
@@
where in the last inequality we further take expectation w.r.t. device $M$ and combine similar terms.
@@
where $1)$ and $2)$ follows from equation (14) and (29) in Lemma V.6 respectively. In $3)$, we let $q_{0}=1$ and combine similar terms. In $4)$, we substitute in (41). Lastly, in $5)$, we denote
@@
Lemma .12 (Restatement of Lemma V.10).
Under Assumption .1-.4, and assume the device delay $\tau_{m}$ is uniformly upper bounded by $\tau$, we have the following lemma:
@@
Before we give the proof of Lemma V.10, we first provide some useful facts that reveal properties of the delayed parameters.
Recall that $\tilde{\chi}^{t}$ denote the delayed parameters on all devices, and $\chi^{t}$ denote the non-delayed version. Let ${\mathcal{F}}_{1}=\mathbb{E}\left[\mbox{$\left\lVert\chi^{t+1}-\chi^{t}\right\rVert$}^{2}\right]$, ${\mathcal{F}}_{2}=\mathbb{E}\left[\mbox{$\left\lVert\tilde{\chi}^{t}-\chi^{t}\right\rVert$}^{2}\right]$.
@@
$\overset{3)}{=}$ $-\eta_{0}(1-P)\left(\frac{1}{4}-\frac{\eta_{0}L(B+8d_{0})}{2B}\right)\mathbb{E}\left[\mbox{$\left\lVert\nabla_{x_{0}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]$
@@
Above, we used Lemma V.9 in step $(1)$, substituted in (45) for ${\mathcal{F}}_{1}$ in step $(2)$, substituted in (46) for ${\mathcal{F}}_{2}$ in step (3), and defined
@@
-C Additional Details on Experiment
In this section, we would like to give a brief introduction of the datasets and model structures and further justify the choice of the experimental designs below. We follow the experimental settings of existing works[chen2020vafl, xie2024improving, castiglia2023flexible] for a fair comparison.
Dataset The choices of datasets cover a large range of tasks, including:
  • MNIST and CIFAR-10: standard image benchmarks in machine learning.
  • ModelNet40: multi-view 3D object classification dataset; each object has 12 views from different angles, which naturally lends itself to feature partitioning across devices in VFL.
  • Amazon Reviews: sentiment analysis task, which we include to evaluate our algorithm in the NLP domain.
Data partition and model selection For MNIST, we use a two-layer CNN model, for VFL’s data partitioning, we split the images by row evenly into 7 sub-images and assign them to 7 devices. For CIFAR-10 we use a four-layer CNN model, and partition each image into 2×2, 4 patches of the same size for 4 devices. For ModelNet40, we use a ResNet-18 model, and partition each object into 12 different camera views and allocate them to 12 devices. For Amazon Reviews, we use a pre-trained BERT [devlin2018bert] model, and split the tokenized data input into 3 paragraphs of the same number of tokens and distributed them across 3 devices. For all four datasets, we use a fully connected model of two linear layers with ReLU activations as the server model.